Find X (Show work) 4,750 - 3,500 = 5,000 *  ($13.44 - x)

Find X (Show work)

4,750 - 3,500 = 5,000 *  ($13.44 - x)
Answer #1

4,750 - 3,500 = 5,000 *  ($13.44 - x)

1250 = 5000*(13.44-X)

1250/5000 = 13.44-X

0.25 = 13.44-X

X = 13.44-0.25

X =13.19